### v1

# Overcoming Uncertain Incompleteness for Robust Multimodal Sequential Diagnosis Prediction via Knowledge Distillation and Random Data Erasing

Abstract
In this paper, we present NECHO v2, a novel framework designed to enhance the predictive accuracy of multimodal sequential patient diagnoses under uncertain missing visit sequences, a common challenge in clinical settings. Firstly, we modify NECHO to handle uncertain modality representation dominance under the imperfect data. Next, we develop a systematic knowledge distillation by employing the modified NECHO as both teacher and student. It encompasses a modality-wise contrastive and hierarchical distillation, transformer representation random distillation, along with other distillations to align representations tightly and effectively. We also utilise random erasing on individual data points within sequences during both training and distillation of teacher to lightly simulate scenario with missing visit information to foster effective knowledge transfer. As a result, NECHO v2 verifies itself by showing superiority in multimodal sequential diagnosis prediction on both balanced and imbalanced incomplete settings on multimodal healthcare data.
Index Terms—  Sequential Diagnosis Prediction, Missing Data, Knowledge Distillation, Multimodal Learning, Data Augmentation
1 Introduction
Predicting future patient diagnoses, a.k.a. sequential diagnosis prediction (SDP), based upon clinical records is crucial for enhancing healthcare decision-making [1, 2, 3, 4, 5]. Recent advances in multimodal learning, which integrate diverse modalities such as clinical notes and demographics, have significantly improved SDP accuracy [4, 5]. Nevertheless, most previous studies assume the availability of all data, which is often impractical due to privacy, equipment failures, and other uncertain factors [5]. Encountering such situation presents a significant challenge to healthcare analytics.
Missing data, a common issue in reality, exacerbates model performance. Basic approaches, such as imputation by mean or excluding incomplete data instances, often fail to preserve true data distribution and result in information loss [6]. Advanced statistical techniques, including Multiple Imputation by Chained Equations (MICE) [7], show better efficacy, but their application in complex multimodal scenarios still remains challenging. Therefore, deep learning approaches such as reconstruction [8, 9, 10], which impute missing features, and knowledge distillation (KD) [11], which transfers teacher’s knowledge on complete data to a student learning with incomplete data [12, 13], have gained popularity.
KD has proven effective in model compression [14, 15, 16] and other applications, such as tackling missing data. Wang et al. [12] modality-specialised teachers transfer knowledge to a multimodal student. MissModal [13] employs geometric multimodal contrastive loss [17] and distribution alignment loss on combination of modality representations. However, no existing methodologies have taken into account the fixed dominance of specific modalities under complete data and the fluctuating modalities importance under incomplete data, leading to sub-optimal performance. Furthermore, there is a research gap in systematically leveraging KD to alleviate the representation discrepancy in teacher-student in missing data contexts.
Meanwhile, data augmentation (DA) is widely used to enhance model performance [18]. Some studies have explored its impact on KD [19]. However, there is still lack of research on effectively applying DA in the presence of incomplete data.
To this end, we propose NECHO v2, not only overcoming the challenges in multimodal SDP with imperfection for the first time but also tackling the aforementioned limitations. First, we modify the original NECHO to manage uncertain modality representation dominance in the presence of missing data. Second, we establish a systematic KD pipeline, including modality-wise contrastive and hierarchical distillation, followed by transformer random representation distillation, penultimate layer distillation, and dual-level logit distillation, to fully transfer teacher’s semantic knowledge from perfect data. We also employ a random data point erasing on visit sequences to simulate missing visit to reduce the data distributional gap and facilitate representation transfer. By doing so, NECHO v2 demonstrates robust superiority under both balanced and imbalanced imperfect data scenarios on MIMIC-III data [20].
2 Methodologies
2.1 Problem Statement
Multimodal EHR Data A clinical record is a time-ordered sequence of visits $V_{1},...,V_{T}$, where $T$ represents the total number of visits for any given patient $P$. Each visit $V_{t}$ at the t-th admission consists of three components: $H_{t}$, demographics; $N_{t}$, a clinical note; and $D_{t}$, a set of diagnosis codes. Specifically, a medical ontology $\mathcal{G}$ is utilised to structure diagnosis codes into three hierarchical levels: unique medical codes, category codes, and disease-typing codes, from leaf to node. The input and output is unique medical codes and category codes, respectively.
Missing Visit Sequences To simulate real-world missingness, we randomly discard aforementioned components in each visit sequence, creating an $m$-modal dataset with $2^{m}$ missing patterns. Missing probabilities are balanced or imbalanced across modalities and kept the same during both training and inference phases.
Fig. 1: The Visualisation of Our Proposed Framework, NECHO v2.
Sequential Diagnosis Prediction Given a patient’s multimodal clinical records for the past $T$ visits under incompleteness, the objective is to predict diagnoses codes that will appear in the $(T+1)$-th visit.
2.2 NECHO v2
In this section, we present the KD-based NECHO v2 framework. For a comprehensive flow detailing the process from input to prediction, refer to the original NECHO [5].
2.2.1 Modification of NECHO
NECHO [5] achieves state-of-the-art performance in SDP by integrating demographics, clinical notes, and medical codes using a medical code-centric framework with bi-modal contrastive loss and a centric multimodal adaptation gate (CMAG) for alignment and fusion. Each modality-specific encoder predicts at the parental level of target medical codes (disease-typing codes) to enhance training.
However, it confronts two issues: 1) under-performance in incomplete data despite outstanding performance in complete data and 2) adoption of a pre-trained BioWord2Vec [21], limiting performance on non-MIMIC-III datasets [20]. To mitigate these concerns, we modify by: 1) replacing the demo $\rightarrow$ code cross-modal transformer (CMT) with a demo $\rightarrow$ note CMT to relieve bias from medical codes and 2) utilisation of clinical TinyBERT [22] as a text encoder to potentially facilitate adaptability to other datasets.
2.2.2 Systematic Knowledge Distillation Framework
Teacher-Student Network Configuration In our KD pipeline, we adopt the modified NECHO as both teacher and student. In addition, the teacher leverages CMAG [5] to consider modality representation dominance when learning with the full data, while the student adopts MAG [23] that adjusts significant representations flexibly, considering fluctuating dominant feature under missing conditions. We avoid using the original NECHO as the teacher to reduce architectural heterogeneity, thereby fostering the KD [24]. We adopt offline distillation [11] where the teacher is trained, then frozen during distillation. Additionally, the teacher is absent during student’s inference.
Modality-wise Contrastive and Hierarchical Distillation We begin our KD process by distilling modality-wise representations from the teacher to the student, using contrastive learning [25] and L2 distance measures (Mean Squared Error, MSE). First, contrastive learning identifies and amplifies both similarities and discrepancies between the representations [25, 17]. When utilised in KD, it encourages the student’s representations to be similar to those of the teacher’s for corresponding samples, while also distinguishing between different samples. MSE further tightens this alignment, reducing deviations and promoting consistency.
Unlike previous method [13], we explicitly distill modality-specific semantic distributions. We utilise a contrastive loss [25] with symmetrical losses to promote stable and effective distillation in a modality-wise fashion. Let teacher and student representations with the same data as positive sample pairs ($\hat{H}_{t}^{m,i}$, $\hat{G}_{t}^{m,i}$), with $m\in(H,N,D)$, and teacher and student are $H$ and $G$, respectively. Then, modality-wise contrastive distillation $\mathcal{L}_{\text{MWCD}}$ is:
$\mathcal{L}_{\text{MWCD}}^{\text{H$\rightarrow$G},m}=-\log\frac{\exp(\langle\hat{H}_{t}^{m,i},\hat{G}_{t}^{m,i}\rangle/\tau)}{\sum_{k=1}^{N}\exp(\langle\hat{H}_{t}^{m,i},\hat{G}_{t}^{m,k}\rangle/\tau)},$ (1)
[…]
where $\langle$,$\rangle$ , ⟩ denotes cosine similarity and the temperature $\tau\in\mathbb{R}^{+}$ is a parameter that controls the distribution concentration and the gradient of the Softmax function. Next, with $\hat{H}_{t}^{m}$ and $\hat{G}_{t}^{m}$, modality-specific teacher and student feature at $t$-th visit, modality-wise hierarchical distillation $\mathcal{L}_{\text{MWHD}}$ via MSE $\|\cdot\|_{2}$ is:
$\mathcal{L}_{\text{MWHD}}=\sum_{m\in\{H,N,D\}}\|\hat{H}_{t}^{m}-\hat{G}_{t}^{m}\|_{2}.$ (4)
Accordingly, the modality-wise contrastive and hierarchical distillation $\mathcal{L}_{\text{MWD}}$ is formulated as the sum of the above two loss terms:
$\mathcal{L}_{\text{MWD}}=\mathcal{L}_{\text{MWCD}}+\mathcal{L}_{\text{MWHD}}.$ (5)
Transformer Representation Random Distillation Previous research have explored intermediate layer distillation (ILD) between transformer layers for compression [15, 26, 22]. Meanwhile, NECHO has cross-modal (CMT) and self-attention transformer encoders (SAT) to align and merge inter- and intra-modality representations. Considering its multiple transformers, layer-wise distillation is computationally expensive. Hence, we distill teacher’s randomly selected final transformer features, reducing computational burden and avoiding overfitting.
Denote representations from two CMTs as $R^{M,H\to N}_{t}$ and $R^{M,D\to N}_{t}$, and those from three SATs are $S^{M,H\to N}_{t}$, $S^{M,D\to N}_{t}$, and $S^{M,D}_{t}$, where $M$ is either teacher or student. Then, for the randomly selected transformer representations, the proposed distillation ($\mathcal{L}_{\text{TR2D}}$) using MSE for both CMT ($\mathcal{L}_{\text{CMTD}}$) and SAT ($\mathcal{L}_{\text{SATD}}$) are:
$\mathcal{L}_{\text{TR2D}}=\mathcal{L}_{\text{CMTD}}+\mathcal{L}_{\text{SATD}},$ (6)
$\text{where}\,\mathcal{L}_{\text{CMTD}}=\sum_{m\in\{H,D\}}\|R^{\text{teacher},m\to N}_{t}-R^{\text{student},m\to N}_{t}\|_{2},$ (7)
$\mathcal{L}_{\text{SATD}}=\sum_{m\in\{H,D\}}\|S^{\text{teacher},m\to N}_{t}-S^{\text{student},m\to N}_{t}\|_{2}+$ (8)
$\|S^{\text{teacher},D}_{t}-S^{\text{student},D}_{t}\|_{2}.$
MAG Distillation To ensure the student model further mimics the teacher, we introduce MAG (penultimate layer) distillation. Its importance is also highlighted due to its rich, informative features [27]. MAG representations from teacher and student be $\text{MAG}^{\text{teacher}}_{t}$ and $\text{MAG}^{\text{student}}_{t}$, respectively. Then, the regarding loss is:
$\mathcal{L}_{\text{MAGD}}=\|\text{MAG}^{\text{teacher}}_{t}-\text{MAG}^{\text{student}}_{t}\|_{2}.$ (9)
Dual Logit Distillation NECHO predicts target codes, as well as parental-level codes (disease typing codes) at the modality-specific encoders. Accordingly, we transfer both teacher predictions to the corresponding student predictions. Previous work [16] argues that MSE outperforms Kullback-Leibler (KL) divergence for logit distillation, without requiring hyper-parameter tuning. Hence, MSE is applied to both distillations.
The final prediction and modality-specific parental-level prediction are $\hat{y}^{M}_{t+1}$ and $\hat{o}^{M,m}_{t+1}$. Then, the dual logit distillation loss $\mathcal{L}_{\text{DualLD}}$ is written as:
$\mathcal{L}_{\text{DualLD}}=\mathcal{L}_{\text{LD}}+\mathcal{L}_{\text{hrchyLD}},$ (10)
$\text{where}\,\mathcal{L}_{\text{LD}}=\|\hat{y}^{\text{teacher}}_{t+1}-\hat{y}^{\text{student}}_{t+1}\|_{2},$ (11)
$\mathcal{L}_{\text{hrchyLD}}=\sum_{m\in\{H,N,D\}}\|\hat{o}^{\text{teacher},m}_{t+1}-\hat{o}^{\text{student},m}_{t+1}\|_{2}$ (12)
where $\mathcal{L}_{\text{LD}}$ and $\mathcal{L}_{\text{hrchyLD}}$ are final logit distillation and modality-specific hierarchical logit distillation, respectively.
Model Optimisation The student model is also optimised using a pair of task loss $\mathcal{L}_{\text{DualCE}}$ (CE stands for Cross Entropy), which consists of two components: one for the target level $\mathcal{L}_{\text{CE}}$ and another for the parental level $\mathcal{L}_{\text{hrchyCE}}$, in accordance with NECHO.
By integrating the task losses with the above distillation losses with the corresponding constant $\lambda$, the full optimisation objective is formulated as:
$\mathcal{L}_{\text{total}}=\lambda_{\text{MWD}}\mathcal{L}_{\text{MWD}}+\lambda_{\text{TR2D}}\mathcal{L}_{\text{TR2D}}+$ (13)
$\lambda_{\text{MAGD}}\mathcal{L}_{\text{MAGD}}+\lambda_{\text{DualLD}}\mathcal{L}_{\text{DualLD}}+\lambda_{\text{DualCE}}\mathcal{L}_{\text{DualCE}}.$
2.2.3 Single Data Point Erasing Guided Curriculum Learning
It is desirable to acquire a teacher that transfers better representations to a student learning with uncertain missing data. Prior studies show that large discrepancies in data distribution between teacher and student can hinder KD [19]. Therefore, we propose random single-point data erasing [28] to both training and distillation of the teacher, a minimalistic approach to mimic missing sequences and alleviate the data distribution gap, thereby further improving KD and the performance of resultant student. Note that, DA is not applied to the student during distillation.
Firstly, the teacher is trained using data erasing guided curriculum learning [29], starting with easier samples and gradually progressing to more difficult ones. Each modality is assigned a missing probability of 0.0 or 0.1 with equal probability until specific epoch $e_{1}$, after which the probability of 0.2 is added. Next, during the distillation, the unified teacher trained in the previous manner, complete data representations are transferred until epoch $e_{2}$, after which training continues with either no missing data or a 0.1 missing ratio.
3 Experiments
3.1 Experimental Setup
Dataset and Pre-processing We evaluate on MIMIC-III data [20], following the pre-processing steps from previous works [2, 5] but with a more rigorous patient selection criteria by removing records 1) with a length of stay of zero or negative and 2) who died within 30 days post-discharge. We also leverage only discharge summaries for clinical notes. Detailed statistics upon pre-processing are in Table 1.
To handle missing data, we assign a value beyond the existing range for erased or missing points in demographics and codes. For instance, if the total number of medical codes is 3882, the missing value is assigned as 3883. We also replace missing tokens in notes with the UNK token.
Criteria MIMIC-III Count General # of Patients 5551 # of Visits 14568 # of Unique Codes 3882 # of Category Codes 126 # of Typing Codes 17 Visit Avg / Max # Visit per Patient 2.62 / 33 Avg / Max # Unique Codes per Visit 13.29 / 39 Avg / Max # Category Codes per Visit 11.46 / 34 Avg / Max # Typing Codes per Visit 6.71 / 15
Table 1: Statistics of MIMIC-III Data After Pre-processing.
Training and Evaluation Details We mostly follow the implementation details from previous study [5]. We set the hidden dimension to 128 and the dropout rate to 0.1. The transformer encoders have 4 heads and 3 layers. We set the temperature $T$ to 0.1 and the alpha $\alpha$ to 0.25 for the contrastive distillation. The coefficients for loss terms are set to 1, except for the $\mathcal{L}_{\text{hrchyCE}}$ which is 0.1.
Optimisation is performed via AdamW [30], with a constant learning rate of 2e-5 for the parameters of clinical TinyBERT and 1e-4 for all other parameters. We train with a batch size of 4 for up to 100 epochs, stopping early if no improvement in validation set for 5 consecutive epochs. For curriculum learning, $e_{1}$ and $e_{2}$ are set to 5 and 10, respectively.
Criteria Models (0.2, 0.2, 0.2) (0.5, 0.5, 0.5) (0.8, 0.8, 0.8) (0.2, 0.2, 0.5) (0.5, 0.2, 0.8) (0.8, 0.2, 0.2) (0.2, 0.8, 0.2) top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 Joint MulT [23] 35.52 51.83 33.77 50.31 30.27 46.74 34.23 50.74 32.71 49.41 36.01 52.78 33.82 50.06 NECHO (original) [5] 35.99 52.99 33.17 49.34 28.96 45.77 34.79 51.51 32.69 50.29 36.96 53.46 33.81 50.34 NECHO (modified, teacher) 36.26 52.72 31.37 47.81 28.86 45.86 34.37 50.93 31.63 49.19 36.85 53.21 34.11 50.40 NECHO (modified, student) 35.96 52.98 33.24 50.73 28.64 46.06 35.29 51.98 33.28 50.68 35.65 52.24 33.28 49.88 KD UnimodalKD [12] 35.45 53.19 32.86 50.25 29.28 46.06 34.18 51.92 33.00 50.61 35.41 52.94 34.32 51.53 MissModal [13] 35.85 52.80 33.41 50.37 30.00 46.84 34.73 51.93 33.17 50.24 35.80 52.43 33.68 51.25 NECHO v2 (ours) 37.02 54.26 34.69 51.13 30.57 47.34 35.3 52.49 34.24 51.01 37.41 53.84 34.73 51.65
Table 2: Experimental Results on Multimodal SDP with Uncertain Missingness on MIMIC-III Data. Missing ratios for each modality are ordered as: Demographics, Clinical Notes, and Medical Codes. Best results are in boldface and the second-best results are underlined.
NECHO v2 is evaluated against joint learning methods (MulT [23] and three NECHO variations: original, teacher, and student [5]) and KD methods (UnimodalKD [12] and MissModal [13]). KD methods use the same teacher and student for fair comparison. Evaluation uses top-$k$ accuracy with $k$ values of 10 and 20, following [1, 3, 5]. Experiments are implemented using PyTorch [31] and trained on a single NVIDIA RTX A6000.
3.2 Experimental Results
3.2.1 Main Results
As shown in Table 2, NECHO v2 demonstrates remarkable performance across various missing scenarios on MIMIC-III dataset. Specifically, it outperforms the original NECHO by 1.52%, its teacher by 3.32%, its student by 1.45%, and UnimodalKD by 1.83% in top-10 accuracy at the balanced missingness of 0.5. Similar trends are observed in other settings. On the other hand, NECHO performs well when medical codes are mostly present (0.2) but predicts poorly in scenarios where codes are highly missing (0.8). This highlights the robust and superior performance of NECHO v2.
This performance gain is attributed to: 1) modifying NECHO for both teacher and student to address varying modality significance, 2) implementing systematic KD, including modality-wise contrastive and hierarchical distillation and transformer representation random distillation and others, to comprehensively mimic teacher at various representation levels, and 3) simulating random visit information by single point data erasing to minimise data distribution gaps and improve KD. These enables student to imitate the teacher in varied incompleteness, ensuring robust performance effectively.
3.2.2 Ablation Studies
Criteria Components (0.2, 0.2, 0.2) (0.5, 0.2, 0.8) top-10 top-20 top-10 top-20 KD w/o $\mathcal{L}_{\text{MWCD}}$ 37.10 54.10 34.37 50.79 w/o $\mathcal{L}_{\text{TR2D}}$ 36.9 53.95 33.15 50.58 w/o $\mathcal{L}_{\text{MAGD}}$ 36.01 53.27 32.91 49.87 w/o $\mathcal{L}_{\text{hrchyLD}}$ 35.58 52.85 34.25 50.96 DA only distillation 36.42 53.32 34.05 50.83 only teacher training 37.28 53.65 32.71 49.72 not at all 36.43 53.68 33.54 50.93 NECHO v2 Full 37.02 54.26 34.24 51.01
Table 3: Ablation Studies on MIMIC-III Data.
To evaluate our proposed components, we conduct ablation studies on MIMIC-III data, as detailed in Table 3. We report two scenarios: a balanced missing ratio of 0.2, and an imbalanced ratios of (0.5, 0.8, 0.2), representing two extremes where medical code representation is either highly dominant or minimal.
We first assess the effectiveness of KD. While $\mathcal{L}_{\text{MWCD}}$ occasionally performs better without it, its consistent use generally enhances performance. The absence of $\mathcal{L}_{\text{TR2D}}$ and $\mathcal{L}_{\text{MAGD}}$ during distillation significantly deteriorates the performance, highlighting the importance of intermediate supervision. Additionally, $\mathcal{L}_{\text{hrchyLD}}$ is beneficial, all of which emphasise systematic KD’s importance to align student’s semantic knowledge to that of the teacher. We also evaluate the efficacy of DA in three scenarios: only distillation, only teacher training, and not at all. Overall performance significantly improves, highlighting the significance of the proposed random single data point erasing under missing visit information.
3.2.3 Comparative Studies
Criteria Components (0.2, 0.2, 0.2) (0.5, 0.2, 0.8) top-10 top-20 top-10 top-20 Pairing Original $\to$ Original 36.79 53.11 32.68 49.53 Original $\to$ Modified w/ MAG 36.44 52.96 34.66 50.99 $\mathcal{L}_{\text{TR2D}}$ Not Random 36.91 53.56 32.83 50.40 NECHO v2 Proposed 37.02 54.26 34.24 51.01
Table 4: Comparative Studies on MIMIC-III Data.
As shown in Table 4, we compare NECHO v2 with various combinations of teacher and student (original to original, original to modified, adopted MAG) and transformer full distillation under the same settings with the ablation studies. Our proposed methodologies achieves the best overall performance, highlighting the importance of: 1) carefully pairing teacher and student to address varying representation dominance and minimise architectural heterogeneity, and 2) introducing randomness into KD to prevent overfitting and foster effective knowledge transfer.
4 Conclusion
We tackle uncertain missing sequences for robust multimodal SDP with NECHO v2. Firstly, we modify NECHO to dynamically adjust dominant representations under varying missingness. Next, we design a systematic KD pipeline, adopting the modified NECHO and leveraging multiple distillations including modality-wise contrastive and hierarchical distillation, to comprehensively learn teacher representations from complete data. We also randomly erase single data points per visit to reduce the data distribution discrepancy and enhance KD process. Extensive experiments show the efficacy of NECHO v2 over the existing methodologies. The code will be released upon publication.

### v2

@@
 Abstract
-In this paper, we present NECHO v2, a novel framework designed to enhance the predictive accuracy of multimodal sequential patient diagnoses under uncertain missing visit sequences, a common challenge in clinical settings.
-Firstly, we modify NECHO to handle uncertain modality representation dominance under the imperfect data.
-Next, we develop a systematic knowledge distillation by employing the modified NECHO as both teacher and student.
-It encompasses a modality-wise contrastive and hierarchical distillation, transformer representation random distillation, along with other distillations to align representations tightly and effectively.
-We also utilise random erasing on individual data points within sequences during both training and distillation of teacher to lightly simulate scenario with missing visit information to foster effective knowledge transfer.
-As a result, NECHO v2 verifies itself by showing superiority in multimodal sequential diagnosis prediction on both balanced and imbalanced incomplete settings on multimodal healthcare data.
-Index Terms—  Sequential Diagnosis Prediction, Missing Data, Knowledge Distillation, Multimodal Learning, Data Augmentation
+In this paper, we present NECHO v2, a novel framework designed to enhance the predictive accuracy of multimodal sequential patient diagnoses under uncertain missing visit sequences, a common challenge in real clinical settings.
+Firstly, we modify NECHO, designed in a diagnosis code-centric fashion, to handle uncertain modality representation dominance under the imperfect data.
+Secondly, we develop a systematic knowledge distillation by employing the modified NECHO as both teacher and student.
+It encompasses a modality-wise contrastive and hierarchical distillation, transformer representation random distillation, along with other distillations to align representations between teacher and student tightly and effectively.
+We also utilise random erasing on individual data points within sequences during both training and distillation of the teacher to lightly simulate scenario with missing visit information, thereby fostering effective knowledge transfer.
+As a result, NECHO v2 verifies itself by showing robust superiority in multimodal sequential diagnosis prediction under both balanced and imbalanced incomplete settings on multimodal healthcare data.
+Index Terms—  Sequential Diagnosis Prediction, Missing Data, Knowledge Distillation, Multimodal Learning, Data Augmentation.
 1 Introduction
@@
 Nevertheless, most previous studies assume the availability of all data, which is often impractical due to privacy, equipment failures, and other uncertain factors [5].
-Encountering such situation presents a significant challenge to healthcare analytics.
+Encountering such situation presents a formidable challenge to healthcare analytics.
 Missing data, a common issue in reality, exacerbates model performance.
@@
 Wang et al.
-[12] modality-specialised teachers transfer knowledge to a multimodal student.
-MissModal [13] employs geometric multimodal contrastive loss [17] and distribution alignment loss on combination of modality representations.
+[12] employs modality-specialised teachers that transfer knowledge to a multimodal student.
+MissModal [13] employs geometric multimodal contrastive loss [17] and distribution alignment loss on a combination of modality representations.
 However, no existing methodologies have taken into account the fixed dominance of specific modalities under complete data and the fluctuating modalities importance under incomplete data, leading to sub-optimal performance.
-Furthermore, there is a research gap in systematically leveraging KD to alleviate the representation discrepancy in teacher-student in missing data contexts.
+Furthermore, there is a research gap in systematically leveraging KD to alleviate the representation discrepancy in teacher-student under missing data.
 Meanwhile, data augmentation (DA) is widely used to enhance model performance [18].
 Some studies have explored its impact on KD [19].
-However, there is still lack of research on effectively applying DA in the presence of incomplete data.
-To this end, we propose NECHO v2, not only overcoming the challenges in multimodal SDP with imperfection for the first time but also tackling the aforementioned limitations.
+However, there is still lack of research on effectively applying DA during KD process in the presence of incomplete data.
+To this end, we propose NECHO v2, not only overcoming the challenges in multimodal SDP with imperfect data for the first time but also tackling the aforementioned limitations.
 First, we modify the original NECHO to manage uncertain modality representation dominance in the presence of missing data.
-Second, we establish a systematic KD pipeline, including modality-wise contrastive and hierarchical distillation, followed by transformer random representation distillation, penultimate layer distillation, and dual-level logit distillation, to fully transfer teacher’s semantic knowledge from perfect data.
+Second, we establish a systematic KD pipeline, including modality-wise contrastive and hierarchical distillation, followed by transformer random representation distillation, MAG distillation, and dual-level logit distillation, to fully transfer the teacher’s semantic knowledge acquired from perfect data.
 We also employ a random data point erasing on visit sequences to simulate missing visit to reduce the data distributional gap and facilitate representation transfer.
-By doing so, NECHO v2 demonstrates robust superiority under both balanced and imbalanced imperfect data scenarios on MIMIC-III data [20].
+By doing so, NECHO v2 shows consistent predominance under both balanced and imbalanced imperfect data scenarios on MIMIC-III data [20].
 2 Methodologies
@@
 Each visit $V_{t}$ at the $t$-th admission consists of three components: $H_{t}$, demographics; $N_{t}$, a clinical note; and $D_{t}$, a set of diagnosis codes.
-Specifically, a medical ontology $\mathcal{G}$ is utilised to structure diagnosis codes into three hierarchical levels: unique medical codes, category codes, and disease-typing codes, from leaf to node.
-The input and output is unique medical codes and category codes, respectively.
-Missing Visit Sequences To simulate real-world missingness, we randomly discard aforementioned components in each visit sequence, creating an $m$-modal dataset with $2^{m}$ missing patterns.
+Specifically, a medical ontology $\mathcal{G}$ is utilised to structure diagnosis codes into three hierarchical levels: unique codes, category codes, and disease-typing codes, from leaf to node.
+Input and output is unique codes and category codes, respectively.
+Missing Visit Sequences To simulate real-world missingness, we randomly discard aforementioned components in each visit sequence, creating an $m$-modal dataset with $12^{m}-$ 1 missing patterns.
 Missing probabilities are balanced or imbalanced across modalities and kept the same during both training and inference phases.
-Fig. 1: The Visualisation of Our Proposed Framework, NECHO v2.
+Fig. 1: A Visualisation of Our Proposed Framework, NECHO v2.
 Sequential Diagnosis Prediction Given a patient’s multimodal clinical records for the past $T$ visits under incompleteness, the objective is to predict diagnoses codes that will appear in the $(T+1)$-th visit.
@@
 2.2.1 Modification of NECHO
-NECHO [5] achieves state-of-the-art performance in SDP by integrating demographics, clinical notes, and medical codes using a medical code-centric framework with bi-modal contrastive loss and a centric multimodal adaptation gate (CMAG) for alignment and fusion.
-Each modality-specific encoder predicts at the parental level of target medical codes (disease-typing codes) to enhance training.
+NECHO [5] achieves state-of-the-art performance in SDP by integrating demographics, clinical notes, and diagnosis codes using a diagnosis code-centric framework with bi-modal contrastive loss and a centric multimodal adaptation gate (CMAG) for alignment and fusion.
+Each modality-specific encoder predicts at the parental level of target diagnosis codes (disease-typing codes) to enhance training.
 However, it confronts two issues: 1) under-performance in incomplete data despite outstanding performance in complete data and 2) adoption of a pre-trained BioWord2Vec [21], limiting performance on non-MIMIC-III datasets [20].
-To mitigate these concerns, we modify by: 1) replacing the demo $\rightarrow$ code cross-modal transformer (CMT) with a demo $\rightarrow$ note CMT to relieve bias from medical codes and 2) utilisation of clinical TinyBERT [22] as a text encoder to potentially facilitate adaptability to other datasets.
+To mitigate these concerns, we modify by: 1) replacing the demo $\rightarrow$ code cross-modal transformer (CMT) with a demo $\rightarrow$ note CMT to relieve bias from diagnosis codes and 2) utilisation of clinical TinyBERT [22] as a note encoder to potentially facilitate adaptability to other datasets.
 2.2.2 Systematic Knowledge Distillation Framework
 Teacher-Student Network Configuration In our KD pipeline, we adopt the modified NECHO as both teacher and student.
-In addition, the teacher leverages CMAG [5] to consider modality representation dominance when learning with the full data, while the student adopts MAG [23] that adjusts significant representations flexibly, considering fluctuating dominant feature under missing conditions.
+In addition, the teacher leverages CMAG [5] to consider modality representation dominance when learning with the full data, whilst the student adopts MAG [23] that adjusts significant representations flexibly, considering fluctuating dominant feature under missing conditions.
 We avoid using the original NECHO as the teacher to reduce architectural heterogeneity, thereby fostering the KD [24].
 We adopt offline distillation [11] where the teacher is trained, then frozen during distillation.
-Additionally, the teacher is absent during student’s inference.
+Additionally, the teacher is absent during the student’s inference.
 Modality-wise Contrastive and Hierarchical Distillation We begin our KD process by distilling modality-wise representations from the teacher to the student, using contrastive learning [25] and L2 distance measures (Mean Squared Error, MSE).
 First, contrastive learning identifies and amplifies both similarities and discrepancies between the representations [25, 17].
-When utilised in KD, it encourages the student’s representations to be similar to those of the teacher’s for corresponding samples, while also distinguishing between different samples.
+When utilised in KD, it encourages the student’s representations to be similar to those of the teacher’s for corresponding samples, whilst also distinguishing between different samples.
 MSE further tightens this alignment, reducing deviations and promoting consistency.
@@
 We utilise a contrastive loss [25] with symmetrical losses to promote stable and effective distillation in a modality-wise fashion.
-Let teacher and student representations with the same data as positive sample pairs ($\hat{H}_{t}^{m,i}$, $\hat{G}_{t}^{m,i}$), with $m\in(H,N,D)$, and teacher and student are $H$ and $G$, respectively.
+Let teacher and student representations with the same data as positive sample pairs ($\hat{R}_{t}^{\text{teacher},m,i}$, $\hat{R}_{t}^{\text{student},m,i}$), with $m\in(H,N,D)$, respectively.
 Then, modality-wise contrastive distillation $\mathcal{L}_{\text{MWCD}}$ is:
-$\mathcal{L}_{\text{MWCD}}^{\text{H$\rightarrow$G},m}=-\log\frac{\exp(\langle\hat{H}_{t}^{m,i},\hat{G}_{t}^{m,i}\rangle/\tau)}{\sum_{k=1}^{N}\exp(\langle\hat{H}_{t}^{m,i},\hat{G}_{t}^{m,k}\rangle/\tau)},$ (1)
-[…]
-[…]
+$\mathcal{L}_{\text{MWCD}}^{\text{teacher $\rightarrow$ student},m}=-\log\frac{\exp(\langle\hat{R}_{t}^{\text{teacher},m,i},\hat{R}_{t}^{\text{student},m,i}\rangle/\tau)}{\sum_{k=1}^{N}\exp(\langle\hat{R}_{t}^{\text{teacher},m,i},\hat{R}_{t}^{\text{student},m,k}\rangle/\tau)},$ (1)
+$\mathcal{L}_{\text{MWCD}}^{\text{student $\rightarrow$ teacher},m}=-\log\frac{\exp(\langle\hat{R}_{t}^{\text{student},m,i},\hat{R}_{t}^{\text{student},m,i}\rangle/\tau)}{\sum_{k=1}^{N}\exp(\langle\hat{R}_{t}^{\text{student},m,i},\hat{R}_{t}^{\text{student},m,k}\rangle/\tau)},$ (2)
+$\mathcal{L}_{\text{MWCD}}=\sum_{m\{H,N,D\}}\{\alpha\mathcal{L}_{\text{MWCD}}^{\text{\text{teacher}, $\rightarrow$ student},m}+(1-\alpha)\mathcal{L}_{\text{MWCD}}^{\text{student $\rightarrow$ \text{teacher}, },m}\}$ (3)
 where $\langle$,$\rangle$ , ⟩ denotes cosine similarity and the temperature $\tau\in\mathbb{R}^{+}$ is a parameter that controls the distribution concentration and the gradient of the Softmax function.
-Next, with $\hat{H}_{t}^{m}$ and $\hat{G}_{t}^{m}$, modality-specific teacher and student feature at $t$-th visit, modality-wise hierarchical distillation $\mathcal{L}_{\text{MWHD}}$ via MSE $\|\cdot\|_{2}$ is:
-$\mathcal{L}_{\text{MWHD}}=\sum_{m\in\{H,N,D\}}\|\hat{H}_{t}^{m}-\hat{G}_{t}^{m}\|_{2}.$ (4)
+Next, with $\hat{R}_{t}^{\text{teacher},m}$ and $\hat{R}_{t}^{\text{student},m}$, modality-specific teacher and student feature at $t$-th visit, modality-wise hierarchical distillation $\mathcal{L}_{\text{MWHD}}$ via MSE $\|\cdot\|_{2}$ is:
+$\mathcal{L}_{\text{MWHD}}=\sum_{m\in\{H,N,D\}}\|\hat{R}_{t}^{\text{teacher},m}-\hat{R}_{t}^{\text{student},m}\|_{2}.$ (4)
 Accordingly, the modality-wise contrastive and hierarchical distillation $\mathcal{L}_{\text{MWD}}$ is formulated as the sum of the above two loss terms:
@@
 Hence, we distill teacher’s randomly selected final transformer features, reducing computational burden and avoiding overfitting.
-Denote representations from two CMTs as $R^{M,H\to N}_{t}$ and $R^{M,D\to N}_{t}$, and those from three SATs are $S^{M,H\to N}_{t}$, $S^{M,D\to N}_{t}$, and $S^{M,D}_{t}$, where $M$ is either teacher or student.
+Denote representations from two CMTs as $C^{M,H\to N}_{t}$ and $C^{M,D\to N}_{t}$, and those from three SATs are $S^{M,H\to N}_{t}$, $S^{M,D\to N}_{t}$, and $S^{M,D}_{t}$, where $M$ is either teacher or student.
 Then, for the randomly selected transformer representations, the proposed distillation ($\mathcal{L}_{\text{TR2D}}$) using MSE for both CMT ($\mathcal{L}_{\text{CMTD}}$) and SAT ($\mathcal{L}_{\text{SATD}}$) are:
 $\mathcal{L}_{\text{TR2D}}=\mathcal{L}_{\text{CMTD}}+\mathcal{L}_{\text{SATD}},$ (6)
-$\text{where}\,\mathcal{L}_{\text{CMTD}}=\sum_{m\in\{H,D\}}\|R^{\text{teacher},m\to N}_{t}-R^{\text{student},m\to N}_{t}\|_{2},$ (7)
+$\text{where}\,\mathcal{L}_{\text{CMTD}}=\sum_{m\in\{H,D\}}\|C^{\text{teacher},m\to N}_{t}-C^{\text{student},m\to N}_{t}\|_{2},$ (7)
 $\mathcal{L}_{\text{SATD}}=\sum_{m\in\{H,D\}}\|S^{\text{teacher},m\to N}_{t}-S^{\text{student},m\to N}_{t}\|_{2}+$ (8)
@@
 Its importance is also highlighted due to its rich, informative features [27].
-MAG representations from teacher and student be $\text{MAG}^{\text{teacher}}_{t}$ and $\text{MAG}^{\text{student}}_{t}$, respectively.
-Then, the regarding loss is:
-$\mathcal{L}_{\text{MAGD}}=\|\text{MAG}^{\text{teacher}}_{t}-\text{MAG}^{\text{student}}_{t}\|_{2}.$ (9)
+MAG representations from teacher and student be $\text{CMAG}_{t}$ and $\text{MAG}_{t}$, respectively.
+The regarding loss is:
+$\mathcal{L}_{\text{MAGD}}=\|\text{CMAG}_{t}-\text{MAG}_{t}\|_{2}.$ (9)
 Dual Logit Distillation NECHO predicts target codes, as well as parental-level codes (disease typing codes) at the modality-specific encoders.
@@
 Model Optimisation The student model is also optimised using a pair of task loss $\mathcal{L}_{\text{DualCE}}$ (CE stands for Cross Entropy), which consists of two components: one for the target level $\mathcal{L}_{\text{CE}}$ and another for the parental level $\mathcal{L}_{\text{hrchyCE}}$, in accordance with NECHO.
-By integrating the task losses with the above distillation losses with the corresponding constant $\lambda$, the full optimisation objective is formulated as:
+By integrating the task losses with the aforementioned distillations losses with each constant $\lambda$, the full optimisation objective is formulated as:
 $\mathcal{L}_{\text{TOTAL}}=\lambda_{\text{MWD}}\mathcal{L}_{\text{MWD}}+\lambda_{\text{TR2D}}\mathcal{L}_{\text{TR2D}}+$ (13)
@@
 2.2.3 Single Data Point Erasing Guided Curriculum Learning
-It is desirable to acquire a teacher that transfers better representations to a student learning with uncertain missing data.
 Prior studies show that large discrepancies in data distribution between teacher and student can hinder KD [19].
@@
 Each modality is assigned a missing probability of 0.0 or 0.1 with equal probability until specific epoch $e_{1}$, after which the probability of 0.2 is added.
-Next, during the distillation, the unified teacher trained in the previous manner, complete data representations are transferred until epoch $e_{2}$, after which training continues with either no missing data or a 0.1 missing ratio.
+Next, during the distillation, the single teacher trained in the previous manner, complete data representations are transferred until epoch $e_{2}$, after which training continues with either no missing data or a 0.1 missing ratio.
+This strategy improves robustness of the teacher against missing data during training and reduces data distribution discrepancies during distillation, leading to an improved representation transfer.
 3 Experiments
 3.1 Experimental Setup
-Dataset and Pre-processing We evaluate on MIMIC-III data [20], following the pre-processing steps from previous works [2, 5] but with a more rigorous patient selection criteria by removing records 1) with a length of stay of zero or negative and 2) who died within 30 days post-discharge.
+Dataset and Pre-processing We evaluate on MIMIC-III data [20], following the pre-processing steps from previous works [2, 5] but with a more rigorous patient selection criteria by removing records of: 1) with a length of stay of non-positive and 2) who died within 30 days post-discharge.
 We also leverage only discharge summaries for clinical notes.
 Detailed statistics upon pre-processing are in Table 1.
-To handle missing data, we assign a value beyond the existing range for erased or missing points in demographics and codes.
-For instance, if the total number of medical codes is 3882, the missing value is assigned as 3883.
+To handle missing data, we assign a value beyond the existing range in demographics and codes.
+For instance, if the total number of diagnosis codes is 3882, the missing value is assigned as 3883.
 We also replace missing tokens in notes with the UNK token.
@@
 For curriculum learning, $e_{1}$ and $e_{2}$ are set to 5 and 10, respectively.
-Criteria Models (0.2, 0.2, 0.2) (0.5, 0.5, 0.5) (0.8, 0.8, 0.8) (0.2, 0.2, 0.5) (0.5, 0.2, 0.8) (0.8, 0.2, 0.2) (0.2, 0.8, 0.2) top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 Joint MulT [23] 35.52 51.83 33.77 50.31 30.27 46.74 34.23 50.74 32.71 49.41 36.01 52.78 33.82 50.06 NECHO (original) [5] 35.99 52.99 33.17 49.34 28.96 45.77 34.79 51.51 32.69 50.29 36.96 53.46 33.81 50.34 NECHO (modified, teacher) 36.26 52.72 31.37 47.81 28.86 45.86 34.37 50.93 31.63 49.19 36.85 53.21 34.11 50.40 NECHO (modified, student) 35.96 52.98 33.24 50.73 28.64 46.06 35.29 51.98 33.28 50.68 35.65 52.24 33.28 49.88 KD UnimodalKD [12] 35.45 53.19 32.86 50.25 29.28 46.06 34.18 51.92 33.00 50.61 35.41 52.94 34.32 51.53 MissModal [13] 35.85 52.80 33.41 50.37 30.00 46.84 34.73 51.93 33.17 50.24 35.80 52.43 33.68 51.25 NECHO v2 (ours) 37.02 54.26 34.69 51.13 30.57 47.34 35.3 52.49 34.24 51.01 37.41 53.84 34.73 51.65
+Criteria Models (0.2, 0.2, 0.2) (0.5, 0.5, 0.5) (0.8, 0.8, 0.8) (0.2, 0.2, 0.5) (0.5, 0.2, 0.8) (0.8, 0.2, 0.2) (0.2, 0.8, 0.2) top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 Joint MulT [23] 35.52 51.83 33.77 50.31 30.27 46.74 34.23 50.74 32.71 49.41 36.01 52.78 33.82 50.06 NECHO (original) [5] 35.99 52.99 33.17 49.34 28.96 45.77 34.79 51.51 32.69 50.29 36.96 53.46 33.81 50.34 NECHO (modified, teacher) 36.26 52.72 31.37 47.81 28.86 45.86 34.37 50.93 31.63 49.19 36.85 53.21 34.11 50.40 NECHO (modified, student) 35.96 52.98 33.24 50.73 28.64 46.06 35.29 51.98 33.28 50.68 35.65 52.24 33.28 49.88 KD UnimodalKD [12] 35.45 53.19 32.86 50.25 29.28 46.06 34.18 51.92 33.00 50.61 35.41 52.94 34.32 51.53 MissModal [13] 35.85 52.80 33.41 50.37 30.00 46.84 34.73 51.93 33.17 50.24 35.80 52.43 33.68 51.25 NECHO v2 (ours) 37.02 54.26 34.69 51.13 30.57 47.34 35.30 52.49 34.24 51.01 37.41 53.84 34.73 51.65
 Table 2: Experimental Results on Multimodal SDP with Uncertain Missingness on MIMIC-III Data.
-Missing ratios for each modality are ordered as: Demographics, Clinical Notes, and Medical Codes.
+Missing ratios for each modality are ordered as: demographics, clinical notes, and diagnosis Codes.
 Best results are in boldface and the second-best results are underlined.
 NECHO v2 is evaluated against joint learning methods (MulT [23] and three NECHO variations: original, teacher, and student [5]) and KD methods (UnimodalKD [12] and MissModal [13]).
-KD methods use the same teacher and student for fair comparison.
+KD methods use the same teacher (or its encoders) and student for fair comparison.
 Evaluation uses top-$k$ accuracy with $k$ values of 10 and 20, following [1, 3, 5].
@@
 As shown in Table 2, NECHO v2 demonstrates remarkable performance across various missing scenarios on MIMIC-III dataset.
-Specifically, it outperforms the original NECHO by 1.52%, its teacher by 3.32%, its student by 1.45%, and UnimodalKD by 1.83% in top-10 accuracy at the balanced missingness of 0.5.
+Specifically, it outperforms MulT by 0.92%, the original NECHO by 1.52%, its teacher by 3.32%, its student by 1.45%, and UnimodalKD by 1.83% in top-10 accuracy at the balanced missingness of 0.5.
 Similar trends are observed in other settings.
-On the other hand, NECHO performs well when medical codes are mostly present (0.2) but predicts poorly in scenarios where codes are highly missing (0.8).
-This highlights the robust and superior performance of NECHO v2.
-This performance gain is attributed to: 1) modifying NECHO for both teacher and student to address varying modality significance, 2) implementing systematic KD, including modality-wise contrastive and hierarchical distillation and transformer representation random distillation and others, to comprehensively mimic teacher at various representation levels, and 3) simulating random visit information by single point data erasing to minimise data distribution gaps and improve KD.
-These enables student to imitate the teacher in varied incompleteness, ensuring robust performance effectively.
+In contrary, NECHO performs well when diagnosis codes are mostly present (0.2) but predicts poorly in scenarios where codes are highly missing (0.8).
+This highlights the consistent supremacy of our NECHO v2.
+This performance gain is attributed to: 1) modifying NECHO for both teacher and student to address varying modality significance, 2) implementing systematic KD, including modality-wise contrastive and hierarchical distillation and transformer representation random distillation, MAG distillation, and dual-level logit distillation, to comprehensively mimic teacher at various representation levels, and 3) simulating random missing visit information by single point data random erasing to minimise data distribution gaps.
+These enables the student to imitate the teacher in varied incompleteness settings, ensuring considerable performance gain effectively.
 3.2.2 Ablation Studies
@@
 To evaluate our proposed components, we conduct ablation studies on MIMIC-III data, as detailed in Table 3.
-We report two scenarios: a balanced missing ratio of 0.2, and an imbalanced ratios of (0.5, 0.8, 0.2), representing two extremes where medical code representation is either highly dominant or minimal.
+We report two scenarios: a balanced missing ratio of 0.2, and an imbalanced ratios of (0.5, 0.8, 0.2), representing two extremes where diagnosis codes representation is either highly dominant or minimal.
 We first assess the effectiveness of KD.
-While $\mathcal{L}_{\text{MWCD}}$ occasionally performs better without it, its consistent use generally enhances performance.
+Whilst $\mathcal{L}_{\text{MWCD}}$ occasionally performs better without it, its consistent use generally enhances performance.
 The absence of $\mathcal{L}_{\text{TR2D}}$ and $\mathcal{L}_{\text{MAGD}}$ during distillation significantly deteriorates the performance, highlighting the importance of intermediate supervision.
-Additionally, $\mathcal{L}_{\text{hrchyLD}}$ is beneficial, all of which emphasise systematic KD’s importance to align student’s semantic knowledge to that of the teacher.
+Additionally, $\mathcal{L}_{\text{hrchyLD}}$ is beneficial in performance.
+The quantitative results validate the importance of all components in systematic KD to align student’s semantic knowledge to that of the teacher.
 We also evaluate the efficacy of DA in three scenarios: only distillation, only teacher training, and not at all.
 Overall performance significantly improves, highlighting the significance of the proposed random single data point erasing under missing visit information.
+This enhances the teacher’s robustness against missingness during training and minimises data distribution discrepancies during distillation, resulting in a student model that is highly resilient to uncertain data incompleteness.
 3.2.3 Comparative Studies
@@
 Table 4: Comparative Studies on MIMIC-III Data.
-As shown in Table 4, we compare NECHO v2 with various combinations of teacher and student (original to original, original to modified, adopted MAG) and transformer full distillation under the same settings with the ablation studies.
+As shown in Table 4, we compare NECHO v2 with various combinations of teacher and student (original to original, original to modified with MAG) and transformer full distillation under the same settings with the ablation studies.
 Our proposed methodologies achieves the best overall performance, highlighting the importance of: 1) carefully pairing teacher and student to address varying representation dominance and minimise architectural heterogeneity, and 2) introducing randomness into KD to prevent overfitting and foster effective knowledge transfer.
 4 Conclusion
-We tackle uncertain missing sequences for robust multimodal SDP with NECHO v2.
-Firstly, we modify NECHO to dynamically adjust dominant representations under varying missingness.
-Next, we design a systematic KD pipeline, adopting the modified NECHO and leveraging multiple distillations including modality-wise contrastive and hierarchical distillation, to comprehensively learn teacher representations from complete data.
-We also randomly erase single data points per visit to reduce the data distribution discrepancy and enhance KD process.
-Extensive experiments show the efficacy of NECHO v2 over the existing methodologies.
-The code will be released upon publication.
+We tackle uncertain missing sequences for robust multimodal SDP with the proposed NECHO v2.
+Extensive experiments show the stable superiority of NECHO v2 over the existing methods.
+The code will be released soon.

### v3

@@
-Overcoming Uncertain Incompleteness for Robust Multimodal Sequential Diagnosis Prediction via Knowledge Distillation and Random Data Erasing
+Overcoming Uncertain Incompleteness for Robust Multimodal Sequential Diagnosis Prediction via Curriculum Data Erasing Guided Distillation
 Abstract
@@
 It encompasses a modality-wise contrastive and hierarchical distillation, transformer representation random distillation, along with other distillations to align representations between teacher and student tightly and effectively.
-We also utilise random erasing on individual data points within sequences during both training and distillation of the teacher to lightly simulate scenario with missing visit information, thereby fostering effective knowledge transfer.
+We also propose curriculum learning guided random data erasing within sequences during both training and distillation of the teacher to lightly simulate scenario with missing visit information, thereby fostering effective knowledge transfer.
 As a result, NECHO v2 verifies itself by showing robust superiority in multimodal sequential diagnosis prediction under both balanced and imbalanced incomplete settings on multimodal healthcare data.
@@
 Predicting future patient diagnoses, a.k.a.
-sequential diagnosis prediction (SDP), based upon clinical records is crucial for enhancing healthcare decision-making [1, 2, 3, 4, 5].
+sequential (next visit) diagnosis prediction (SDP), based upon clinical records is crucial for enhancing healthcare decision-making [1, 2, 3, 4].
 Recent advances in multimodal learning, which integrate diverse modalities such as clinical notes and demographics, have significantly improved SDP accuracy [4, 1].
 Nevertheless, most previous studies assume the availability of all data, which is often impractical due to privacy, equipment failures, and other uncertain factors [1].
-Encountering such situation presents a formidable challenge to healthcare analytics.
+Encountering such situations presents a formidable challenge to healthcare analytics.
 Missing data, a common issue in reality, exacerbates model performance.
 Basic approaches, such as imputation by mean or excluding incomplete data instances, often fail to preserve true data distribution and result in information loss [5].
-Advanced statistical techniques, including Multiple Imputation by Chained Equations (MICE) [7], show better efficacy, but their application in complex multimodal scenarios still remains challenging.
+Advanced statistical techniques, including Multivariate Imputation by Chained Equations (MICE) [6], show better efficacy, but their application in complex multimodal scenarios still remains challenging.
 Therefore, deep learning approaches such as reconstruction [7, 8, 9], which impute missing features, and knowledge distillation (KD) [10], which transfers teacher’s knowledge on complete data to a student learning with incomplete data [11, 12], have gained popularity.
@@
 Wang et al.
-[12] employs modality-specialised teachers that transfer knowledge to a multimodal student.
-MissModal [13] employs geometric multimodal contrastive loss [17] and distribution alignment loss on a combination of modality representations.
-However, no existing methodologies have taken into account the fixed dominance of specific modalities under complete data and the fluctuating modalities importance under incomplete data, leading to sub-optimal performance.
-Furthermore, there is a research gap in systematically leveraging KD to alleviate the representation discrepancy in teacher-student under missing data.
-Meanwhile, data augmentation (DA) is widely used to enhance model performance [18].
-Some studies have explored its impact on KD [19].
-However, there is still lack of research on effectively applying DA during KD process in the presence of incomplete data.
+[11] employs modality-specialised teachers that migrate knowledge to a multimodal student.
+MissModal [12] employs geometric multimodal contrastive loss [16] and distribution alignment loss on a combination of modality representations in a self-distillation manner [17].
+However, there is a research gap in systematically leveraging KD to alleviate the representation discrepancy in teacher-student under missing data.
+Furthermore, no existing methodologies have taken into account the fixed dominance of specific modalities under complete data and the fluctuating modalities importance under incomplete data, leading to sub-optimal performance.
+Meanwhile, some studies examine the impact of data augmentation [18] on KD [19, 20]. However, research on applying data augmentation during KD with incomplete data remains under-explored.
 To this end, we propose NECHO v2, not only overcoming the challenges in multimodal SDP with imperfect data for the first time but also tackling the aforementioned limitations.
-First, we modify the original NECHO to manage uncertain modality representation dominance in the presence of missing data.
-Second, we establish a systematic KD pipeline, including modality-wise contrastive and hierarchical distillation, followed by transformer random representation distillation, MAG distillation, and dual-level logit distillation, to fully transfer the teacher’s semantic knowledge acquired from perfect data.
-We also employ a random data point erasing on visit sequences to simulate missing visit to reduce the data distributional gap and facilitate representation transfer.
+First, we modify the original NECHO to manage uncertain modality dominance in the presence of missing data.
+Second, we establish a systematic KD pipeline, including modality-wise contrastive and hierarchical distillation, followed by transformer random representation distillation, MAG distillation, and dual-level logit distillation, to fully transfer the teacher’s semantic knowledge acquired from the perfect data.
+Lastly, we introduce a random data erasing on each visit sequence in a curriculum fashion, simulating missing visit to reduce data distributional gap and facilitate representation propagation.
 By doing so, NECHO v2 shows consistent predominance under both balanced and imbalanced imperfect data scenarios on MIMIC-III data [21].
@@
 2.1 Problem Statement
-Multimodal EHR Data A clinical record is a time-ordered sequence of visits $V_{1},...,V_{T}$, where $T$ represents the total number of visits for any given patient $P$.
-Each visit $V_{t}$ at the $t$-th admission consists of three components: $H_{t}$, demographics; $N_{t}$, a clinical note; and $D_{t}$, a set of diagnosis codes.
+Multimodal EHR Data.
+A clinical record is a time-ordered sequence of visits $V_{1},...,V_{T}$, where $T$ represents the total number of visits for any given patient $P$.
+Each visit $V_{t}$ at $t$-th admission consists of three components: $D$, demographics; $N$, a clinical note; and $C$, a set of diagnosis codes.
 Specifically, a medical ontology $\mathcal{G}$ is utilised to structure diagnosis codes into three hierarchical levels: unique codes, category codes, and disease-typing codes, from leaf to node.
-Input and output is unique codes and category codes, respectively.
-Missing Visit Sequences To simulate real-world missingness, we randomly discard aforementioned components in each visit sequence, creating an $m$-modal dataset with $12^{m}-$ 1 missing patterns.
+Input and target are unique codes and category codes.
+Missing Visit Sequences.
+To simulate real-world missingness, we randomly discard aforementioned components in each visit sequence, creating an $m$-modal dataset with $12^{m}-$ 1 missing patterns.
 Missing probabilities are balanced or imbalanced across modalities and kept the same during both training and inference phases.
 Fig. 1: A Visualisation of Our Proposed Framework, NECHO v2.
-Sequential Diagnosis Prediction Given a patient’s multimodal clinical records for the past $T$ visits under incompleteness, the objective is to predict diagnoses codes that will appear in the $(T+1)$-th visit.
+Sequential Diagnosis Prediction. Given a patient’s incomplete multimodal clinical records for the past $T$ visits, the objective is to predict diagnoses codes that will appear in the $(T+1)$-th visit.
 2.2 NECHO v2
@@
 Each modality-specific encoder predicts at the parental level of target diagnosis codes (disease-typing codes) to enhance training.
-However, it confronts two issues: 1) under-performance in incomplete data despite outstanding performance in complete data and 2) adoption of a pre-trained BioWord2Vec [21], limiting performance on non-MIMIC-III datasets [20].
-To mitigate these concerns, we modify by: 1) replacing the demo $\rightarrow$ code cross-modal transformer (CMT) with a demo $\rightarrow$ note CMT to relieve bias from diagnosis codes and 2) utilisation of clinical TinyBERT [22] as a note encoder to potentially facilitate adaptability to other datasets.
+However, it confronts two issues: 1) under-performance under incomplete data despite outstanding performance under complete data, and 2) adoption of a pre-trained BioWord2Vec [22], limiting the adaptation to missing data.
+To mitigate these concerns, we modify by: 1) replacing the demo $\rightarrow$ code transformers with a demo $\rightarrow$ note transformers to relieve bias from diagnosis codes, and 2) utilisation of clinical TinyBERT [23] as a note encoder to potentially facilitate adaptability to incomplete data.
 2.2.2 Systematic Knowledge Distillation Framework
-Teacher-Student Network Configuration In our KD pipeline, we adopt the modified NECHO as both teacher and student.
-In addition, the teacher leverages CMAG [5] to consider modality representation dominance when learning with the full data, whilst the student adopts MAG [23] that adjusts significant representations flexibly, considering fluctuating dominant feature under missing conditions.
-We avoid using the original NECHO as the teacher to reduce architectural heterogeneity, thereby fostering the KD [24].
-We adopt offline distillation [11] where the teacher is trained, then frozen during distillation.
+Teacher-Student Network Configuration.
+In our KD pipeline, we adopt the modified NECHO as both teacher and student.
+Architecturally, the teacher leverages CMAG [1] to consider modality representation dominance when learning with the full data, whilst the student adopts MAG [4] that adjusts significant representations flexibly, considering fluctuating dominant features under missing conditions.
+Additionally, we avoid using the original NECHO as the teacher to reduce architectural heterogeneity, thereby fostering the KD [24].
+For the distillation process, we adopt offline distillation [10] where the teacher is trained, then frozen during distillation.
 Additionally, the teacher is absent during the student’s inference.
-Modality-wise Contrastive and Hierarchical Distillation We begin our KD process by distilling modality-wise representations from the teacher to the student, using contrastive learning [25] and L2 distance measures (Mean Squared Error, MSE).
-First, contrastive learning identifies and amplifies both similarities and discrepancies between the representations [25, 17].
+Modality-wise Contrastive and Hierarchical Distillation.
+We begin our KD process by distilling modality-wise representations from the teacher to the student, using contrastive learning [25, 26] and L2 distance measures (Mean Squared Error, MSE).
+First, contrastive learning identifies and amplifies both similarities and discrepancies between the representations [25, 26, 16].
 When utilised in KD, it encourages the student’s representations to be similar to those of the teacher’s for corresponding samples, whilst also distinguishing between different samples.
 MSE further tightens this alignment, reducing deviations and promoting consistency.
-Unlike previous method [13], we explicitly distill modality-specific semantic distributions.
-We utilise a contrastive loss [25] with symmetrical losses to promote stable and effective distillation in a modality-wise fashion.
-Let teacher and student representations with the same data as positive sample pairs ($\hat{R}_{t}^{\text{teacher},m,i}$, $\hat{R}_{t}^{\text{student},m,i}$), with $m\in(H,N,D)$, respectively.
-Then, modality-wise contrastive distillation $\mathcal{L}_{\text{MWCD}}$ is:
-$\mathcal{L}_{\text{MWCD}}^{\text{teacher $\rightarrow$ student},m}=-\log\frac{\exp(\langle\hat{R}_{t}^{\text{teacher},m,i},\hat{R}_{t}^{\text{student},m,i}\rangle/\tau)}{\sum_{k=1}^{N}\exp(\langle\hat{R}_{t}^{\text{teacher},m,i},\hat{R}_{t}^{\text{student},m,k}\rangle/\tau)},$ (1)
-$\mathcal{L}_{\text{MWCD}}^{\text{student $\rightarrow$ teacher},m}=-\log\frac{\exp(\langle\hat{R}_{t}^{\text{student},m,i},\hat{R}_{t}^{\text{student},m,i}\rangle/\tau)}{\sum_{k=1}^{N}\exp(\langle\hat{R}_{t}^{\text{student},m,i},\hat{R}_{t}^{\text{student},m,k}\rangle/\tau)},$ (2)
-$\mathcal{L}_{\text{MWCD}}=\sum_{m\{H,N,D\}}\{\alpha\mathcal{L}_{\text{MWCD}}^{\text{\text{teacher}, $\rightarrow$ student},m}+(1-\alpha)\mathcal{L}_{\text{MWCD}}^{\text{student $\rightarrow$ \text{teacher}, },m}\}$ (3)
+Unlike previous methods [11, 12], we explicitly distill modality-specific semantic distributions.
+We utilise a contrastive loss with symmetrical losses to promote stable and effective distillation in a modality-wise fashion.
+Let teacher and student representations with the same data as positive sample pairs ($\hat{R}_{t}^{\text{teacher},m,i}$, $\hat{R}_{t}^{\text{student},m,i}$), with $m\in(D,N,C)$, respectively.
+Then, with weighting parameter $\alpha$ and batch size $K$, the modality-wise contrastive distillation $\mathcal{L}_{\text{MWCD}}$ is as follows:
+$\mathcal{L}_{\text{MWCD}}^{\text{teacher $\rightarrow$ student},m}=-\log\frac{\exp(\langle\hat{R}_{t}^{\text{teacher},m,i},\hat{R}_{t}^{\text{student},m,i}\rangle/\tau)}{\sum_{k=1}^{K}\exp(\langle\hat{R}_{t}^{\text{teacher},m,i},\hat{R}_{t}^{\text{student},m,k}\rangle/\tau)},$ (1)
+$\mathcal{L}_{\text{MWCD}}^{\text{student $\rightarrow$ teacher},m}=-\log\frac{\exp(\langle\hat{R}_{t}^{\text{student},m,i},\hat{R}_{t}^{\text{teacher},m,i}\rangle/\tau)}{\sum_{k=1}^{K}\exp(\langle\hat{R}_{t}^{\text{student},m,i},\hat{R}_{t}^{\text{teacher},m,k}\rangle/\tau)},$ (2)
+$\mathcal{L}_{\text{MWCD}}=\sum_{m\{D,N,C\}}\{\alpha\mathcal{L}_{\text{MWCD}}^{\text{\text{teacher}, $\rightarrow$ student},m}+(1-\alpha)\mathcal{L}_{\text{MWCD}}^{\text{student $\rightarrow$ \text{teacher}, },m}\}$ (3)
 where $\langle$,$\rangle$ , ⟩ denotes cosine similarity and the temperature $\tau\in\mathbb{R}^{+}$ is a parameter that controls the distribution concentration and the gradient of the Softmax function.
 Next, with $\hat{R}_{t}^{\text{teacher},m}$ and $\hat{R}_{t}^{\text{student},m}$, modality-specific teacher and student feature at $t$-th visit, modality-wise hierarchical distillation $\mathcal{L}_{\text{MWHD}}$ via MSE $\|\cdot\|_{2}$ is:
-$\mathcal{L}_{\text{MWHD}}=\sum_{m\in\{H,N,D\}}\|\hat{R}_{t}^{\text{teacher},m}-\hat{R}_{t}^{\text{student},m}\|_{2}.$ (4)
+$\mathcal{L}_{\text{MWHD}}=\sum_{m\in\{D,N,C\}}\|\hat{R}_{t}^{\text{teacher},m}-\hat{R}_{t}^{\text{student},m}\|_{2}.$ (4)
 Accordingly, the modality-wise contrastive and hierarchical distillation $\mathcal{L}_{\text{MWD}}$ is formulated as the sum of the above two loss terms:
 $\mathcal{L}_{\text{MWD}}=\mathcal{L}_{\text{MWCD}}+\mathcal{L}_{\text{MWHD}}.$ (5)
-Transformer Representation Random Distillation Previous research have explored intermediate layer distillation (ILD) between transformer layers for compression [15, 26, 22].
+Transformer Representation Random Distillation.
+Previous research have explored intermediate layer distillation (ILD) between transformer layers for compression [14, 27, 23].
 Meanwhile, NECHO has cross-modal (CMT) and self-attention transformer encoders (SAT) to align and merge inter- and intra-modality representations.
@@
 Hence, we distill teacher’s randomly selected final transformer features, reducing computational burden and avoiding overfitting.
-Denote representations from two CMTs as $C^{M,H\to N}_{t}$ and $C^{M,D\to N}_{t}$, and those from three SATs are $S^{M,H\to N}_{t}$, $S^{M,D\to N}_{t}$, and $S^{M,D}_{t}$, where $M$ is either teacher or student.
+Denote representations from two CMTs as $C^{M,D\to N}_{t}$ and $C^{M,C\to N}_{t}$, and those from three SATs are $S^{M,D\to N}_{t}$, $S^{M,C\to N}_{t}$, and $S^{M,C}_{t}$, where $M$ is either teacher or student.
 Then, for the randomly selected transformer representations, the proposed distillation ($\mathcal{L}_{\text{TR2D}}$) using MSE for both CMT ($\mathcal{L}_{\text{CMTD}}$) and SAT ($\mathcal{L}_{\text{SATD}}$) are:
 $\mathcal{L}_{\text{TR2D}}=\mathcal{L}_{\text{CMTD}}+\mathcal{L}_{\text{SATD}},$ (6)
-$\text{where}\,\mathcal{L}_{\text{CMTD}}=\sum_{m\in\{H,D\}}\|C^{\text{teacher},m\to N}_{t}-C^{\text{student},m\to N}_{t}\|_{2},$ (7)
-$\mathcal{L}_{\text{SATD}}=\sum_{m\in\{H,D\}}\|S^{\text{teacher},m\to N}_{t}-S^{\text{student},m\to N}_{t}\|_{2}+$ (8)
-$\|S^{\text{teacher},D}_{t}-S^{\text{student},D}_{t}\|_{2}.$
-MAG Distillation To ensure the student model further mimics the teacher, we introduce MAG (penultimate layer) distillation.
+$\text{where}\,\mathcal{L}_{\text{CMTD}}=\sum_{m\in\{D,C\}}\|C^{\text{teacher},m\to N}_{t}-C^{\text{student},m\to N}_{t}\|_{2},$ (7)
+$\mathcal{L}_{\text{SATD}}=\sum_{m\in\{D,C\}}\|S^{\text{teacher},m\to N}_{t}-S^{\text{student},m\to N}_{t}\|_{2}+$ (8)
+$\|S^{\text{teacher},C}_{t}-S^{\text{student},C}_{t}\|_{2}.$
+MAG Distillation.
+To ensure the student model further mimics the teacher, we introduce MAG (penultimate layer) distillation.
 Its importance is also highlighted due to its rich, informative features [28].
-MAG representations from teacher and student be $\text{CMAG}_{t}$ and $\text{MAG}_{t}$, respectively.
+Let MAG representations from teacher and student be $\text{CMAG}_{t}$ and $\text{MAG}_{t}$, respectively.
 The regarding loss is:
 $\mathcal{L}_{\text{MAGD}}=\|\text{CMAG}_{t}-\text{MAG}_{t}\|_{2}.$ (9)
-Dual Logit Distillation NECHO predicts target codes, as well as parental-level codes (disease typing codes) at the modality-specific encoders.
+Dual Logit Distillation.
+NECHO predicts target codes, as well as parental-level codes (disease typing codes) at the modality-specific encoders.
 Accordingly, we transfer both teacher predictions to the corresponding student predictions.
@@
 $\text{where}\,\mathcal{L}_{\text{LD}}=\|\hat{y}^{\text{teacher}}_{t+1}-\hat{y}^{\text{student}}_{t+1}\|_{2},$ (11)
-$\mathcal{L}_{\text{hrchyLD}}=\sum_{m\in\{H,N,D\}}\|\hat{o}^{\text{teacher},m}_{t+1}-\hat{o}^{\text{student},m}_{t+1}\|_{2}$ (12)
+$\mathcal{L}_{\text{hrchyLD}}=\sum_{m\in\{D,N,C\}}\|\hat{o}^{\text{teacher},m}_{t+1}-\hat{o}^{\text{student},m}_{t+1}\|_{2}$ (12)
 where $\mathcal{L}_{\text{LD}}$ and $\mathcal{L}_{\text{hrchyLD}}$ are final logit distillation and modality-specific hierarchical logit distillation, respectively.
-Model Optimisation The student model is also optimised using a pair of task loss $\mathcal{L}_{\text{DualCE}}$ (CE stands for Cross Entropy), which consists of two components: one for the target level $\mathcal{L}_{\text{CE}}$ and another for the parental level $\mathcal{L}_{\text{hrchyCE}}$, in accordance with NECHO.
-By integrating the task losses with the aforementioned distillations losses with each constant $\lambda$, the full optimisation objective is formulated as:
+Model Optimisation.
+The student model is also optimised using a pair of task loss $\mathcal{L}_{\text{DualCE}}$ (CE stands for Cross Entropy), which consists of two components: one for the target level $\mathcal{L}_{\text{CE}}$ and the other for the parental level $\mathcal{L}_{\text{hrchyCE}}$, in accordance with NECHO.
+By integrating the task losses with the aforementioned distillation losses with each constant $\lambda$, the full optimisation objective is formulated as:
 $\mathcal{L}_{\text{TOTAL}}=\lambda_{\text{MWD}}\mathcal{L}_{\text{MWD}}+\lambda_{\text{TR2D}}\mathcal{L}_{\text{TR2D}}+$ (13)
 $\lambda_{\text{MAGD}}\mathcal{L}_{\text{MAGD}}+\lambda_{\text{DualLD}}\mathcal{L}_{\text{DualLD}}+\lambda_{\text{DualCE}}\mathcal{L}_{\text{DualCE}}.$
-2.2.3 Single Data Point Erasing Guided Curriculum Learning
-Prior studies show that large discrepancies in data distribution between teacher and student can hinder KD [19].
-Therefore, we propose random single-point data erasing [28] to both training and distillation of the teacher, a minimalistic approach to mimic missing sequences and alleviate the data distribution gap, thereby further improving KD and the performance of resultant student.
-Note that, DA is not applied to the student during distillation.
-Firstly, the teacher is trained using data erasing guided curriculum learning [29], starting with easier samples and gradually progressing to more difficult ones.
+2.2.3 Curriculum Learning Guided Random Data Erasing
+Prior study shows that large discrepancies in data distribution between teacher and student can hinder KD [19].
+Therefore, we propose curriculum learning [29] guided random single-point data erasing [30] to both training and distillation of the teacher.
+It is a minimalistic approach to mimic missing sequences and alleviate the data distribution gap to improve KD.
+Note that, it is not applied to the student during the distillation.
+Firstly, the teacher is trained using curriculum learning guided random data erasing, starting with easier samples and gradually progressing to more difficult ones.
 Each modality is assigned a missing probability of 0.0 or 0.1 with equal probability until specific epoch $e_{1}$, after which the probability of 0.2 is added.
-Next, during the distillation, the single teacher trained in the previous manner, complete data representations are transferred until epoch $e_{2}$, after which training continues with either no missing data or a 0.1 missing ratio.
-This strategy improves robustness of the teacher against missing data during training and reduces data distribution discrepancies during distillation, leading to an improved representation transfer.
+Thereafter, during the distillation, complete data representations from the teacher trained in the previous manner are migrated until epoch $e_{2}$, after which training continues with either no missing data or a 0.1 missing ratio.
+This strategy improves robustness of the teacher against missing data during training and reduces data distribution discrepancies during distillation, leading to an improved representation transmission.
 3 Experiments
 3.1 Experimental Setup
-Dataset and Pre-processing We evaluate on MIMIC-III data [20], following the pre-processing steps from previous works [2, 5] but with a more rigorous patient selection criteria by removing records of: 1) with a length of stay of non-positive and 2) who died within 30 days post-discharge.
+Dataset and Pre-processing.
+We evaluate on MIMIC-III data [21], following pre-processing steps from previous works [3, 1] but with a more rigorous patient selection criteria by removing records of: 1) with a length of stay of non-positive, and 2) who died within 30 days post-discharge.
 We also leverage only discharge summaries for clinical notes.
 Detailed statistics upon pre-processing are in Table 1.
-To handle missing data, we assign a value beyond the existing range in demographics and codes.
-For instance, if the total number of diagnosis codes is 3882, the missing value is assigned as 3883.
-We also replace missing tokens in notes with the UNK token.
-Criteria MIMIC-III Count General # of Patients 5551 # of Visits 14568 # of Unique Codes 3882 # of Category Codes 126 # of Typing Codes 17 Visit Avg / Max # Visit per Patient 2.62 / 33 Avg / Max # Unique Codes per Visit 13.29 / 39 Avg / Max # Category Codes per Visit 11.46 / 34 Avg / Max # Typing Codes per Visit 6.71 / 15
+To handle missing data, we assign a value beyond the existing range in demographics and diagnosis codes.
+For instance, if the total number of codes is 3882, the missing value is assigned as 3883.
+We also replace missing tokens in clinical notes with the UNK token [9].
+Criteria MIMIC-III Count General # of Patients 5551 # of Unique Codes 3882 # of Category Codes 126 # of Typing Codes 17 Visit # of Visits 14568 Avg / Max # Visit per Patient 3.37 / 33 Avg / Max # Unique Codes per Visit 13.29 / 39 Avg / Max # Category Codes per Visit 11.46 / 34 Avg / Max # Typing Codes per Visit 6.71 / 15
 Table 1: Statistics of MIMIC-III Data After Pre-processing.
-Training and Evaluation Details We mostly follow the implementation details from previous study [5].
+Training and Evaluation Details.
+We mostly follow the implementation details from previous study [1].
 We set the hidden dimension to 128 and the dropout rate to 0.1.
@@
 We set the temperature $T$ to 0.1 and the alpha $\alpha$ to 0.25 for the contrastive distillation.
-The coefficients for loss terms are set to 1, except for the $\mathcal{L}_{\text{hrchyCE}}$ which is 0.1.
+The coefficients for loss terms are set to 1, except for the $\mathcal{L}_{\text{TR2D}}$ and $\mathcal{L}_{\text{hrchyCE}}$ which are 0.1.
 Optimisation is performed via AdamW [31], with a constant learning rate of 2e-5 for the parameters of clinical TinyBERT and 1e-4 for all other parameters.
@@
 For curriculum learning, $e_{1}$ and $e_{2}$ are set to 5 and 10, respectively.
-Criteria Models (0.2, 0.2, 0.2) (0.5, 0.5, 0.5) (0.8, 0.8, 0.8) (0.2, 0.2, 0.5) (0.5, 0.2, 0.8) (0.8, 0.2, 0.2) (0.2, 0.8, 0.2) top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 Joint MulT [23] 35.52 51.83 33.77 50.31 30.27 46.74 34.23 50.74 32.71 49.41 36.01 52.78 33.82 50.06 NECHO (original) [5] 35.99 52.99 33.17 49.34 28.96 45.77 34.79 51.51 32.69 50.29 36.96 53.46 33.81 50.34 NECHO (modified, teacher) 36.26 52.72 31.37 47.81 28.86 45.86 34.37 50.93 31.63 49.19 36.85 53.21 34.11 50.40 NECHO (modified, student) 35.96 52.98 33.24 50.73 28.64 46.06 35.29 51.98 33.28 50.68 35.65 52.24 33.28 49.88 KD UnimodalKD [12] 35.45 53.19 32.86 50.25 29.28 46.06 34.18 51.92 33.00 50.61 35.41 52.94 34.32 51.53 MissModal [13] 35.85 52.80 33.41 50.37 30.00 46.84 34.73 51.93 33.17 50.24 35.80 52.43 33.68 51.25 NECHO v2 (ours) 37.02 54.26 34.69 51.13 30.57 47.34 35.30 52.49 34.24 51.01 37.41 53.84 34.73 51.65
+Criteria Models (0.2, 0.2, 0.2) (0.5, 0.5, 0.5) (0.8, 0.8, 0.8) (0.2, 0.2, 0.5) (0.2, 0.8, 0.2) (0.5, 0.2, 0.8) (0.5, 0.8, 0.8) (0.8, 0.2, 0.2) (0.8, 0.2, 0.8) top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 top-10 top-20 Joint MulT [32] 35.52 51.83 33.77 50.31 30.27 46.74 34.23 50.74 33.82 50.06 32.71 49.41 30.39 47.29 36.01 52.78 33.58 50.78 NECHO (Original) [1] 35.99 52.99 33.17 49.34 28.96 45.77 35.02 51.40 33.81 50.34 32.69 50.29 29.36 45.61 36.96 53.46 31.60 48.62 NECHO (Modified for Teacher) 36.26 52.72 31.37 47.81 28.86 45.86 34.37 50.93 34.11 50.40 31.63 49.19 30.08 47.01 36.85 53.21 33.20 50.29 NECHO (Modified for Student) 35.96 52.98 33.24 50.73 28.64 46.06 35.29 51.98 33.28 49.88 33.28 50.68 29.05 46.44 35.65 52.24 31.83 49.25 KD UnimodalKD [11] 35.45 53.19 32.86 50.25 29.28 46.06 34.18 51.92 34.32 51.53 33.00 50.61 29.82 46.71 35.41 52.94 33.08 50.15 MissModal [12] 35.85 52.80 33.41 50.37 30.00 46.84 34.73 51.93 33.68 51.25 33.17 50.24 29.68 47.34 35.80 52.43 32.11 50.62 NECHO v2 (Ours) 37.02 54.26 34.69 51.13 30.57 47.34 35.30 52.49 34.73 51.65 34.24 51.01 30.87 48.07 37.41 53.84 34.71 50.94
 Table 2: Experimental Results on Multimodal SDP with Uncertain Missingness on MIMIC-III Data.
@@
 Best results are in boldface and the second-best results are underlined.
-NECHO v2 is evaluated against joint learning methods (MulT [23] and three NECHO variations: original, teacher, and student [5]) and KD methods (UnimodalKD [12] and MissModal [13]).
+NECHO v2 is evaluated against joint learning methods (MulT [32] and three NECHO [1] variations: original, teacher, and student) and KD methods (UnimodalKD [11] and MissModal [12]).
 KD methods use the same teacher (or its encoders) and student for fair comparison.
-Evaluation uses top-$k$ accuracy with $k$ values of 10 and 20, following [1, 3, 5].
-Experiments are implemented using PyTorch [31] and trained on a single NVIDIA RTX A6000.
+Evaluation uses top-$k$ accuracy with $k$ values of 10 and 20, following [2, 1].
+Experiments are implemented using PyTorch [33] and conducted on a single NVIDIA RTX A6000.
 3.2 Experimental Results
 3.2.1 Main Results
-As shown in Table 2, NECHO v2 demonstrates remarkable performance across various missing scenarios on MIMIC-III dataset.
+As shown in Table 2, NECHO v2 demonstrates remarkable performance across various missingness scenarios on MIMIC-III dataset.
 Specifically, it outperforms MulT by 0.92%, the original NECHO by 1.52%, its teacher by 3.32%, its student by 1.45%, and UnimodalKD by 1.83% in top-10 accuracy at the balanced missingness of 0.5.
 Similar trends are observed in other settings.
-In contrary, NECHO performs well when diagnosis codes are mostly present (0.2) but predicts poorly in scenarios where codes are highly missing (0.8).
-This highlights the consistent supremacy of our NECHO v2.
-This performance gain is attributed to: 1) modifying NECHO for both teacher and student to address varying modality significance, 2) implementing systematic KD, including modality-wise contrastive and hierarchical distillation and transformer representation random distillation, MAG distillation, and dual-level logit distillation, to comprehensively mimic teacher at various representation levels, and 3) simulating random missing visit information by single point data random erasing to minimise data distribution gaps.
+In contrast, NECHO performs well when diagnosis codes are mostly present (0.2) but predicts poorly in scenarios where codes are highly missing (0.8).
+UnimodalKD and Missmodal underperform in most incomplete scenarios, highlighting the need for systematic knowledge distillation that accounts for fluctuating modality dominance under imperfect data.
+This remarkable performance gain of NECHO v2 is attributed to: 1) modifying NECHO to manage varying modality significance under imperfect data, 2) implementing systematic KD, including modality-wise contrastive and hierarchical distillation, to comprehensively mimic teacher at various representation levels, and 3) simulating random missing visit information by curriculum random data erasing to minimise data distribution gaps.
 These enables the student to imitate the teacher in varied incompleteness settings, ensuring considerable performance gain effectively.
 3.2.2 Ablation Studies
-Criteria Components (0.2, 0.2, 0.2) (0.5, 0.2, 0.8) top-10 top-20 top-10 top-20 KD w/o $\mathcal{L}_{\text{MWCD}}$ 37.10 54.10 34.37 50.79 w/o $\mathcal{L}_{\text{TR2D}}$ 36.9 53.95 33.15 50.58 w/o $\mathcal{L}_{\text{MAGD}}$ 36.01 53.27 32.91 49.87 w/o $\mathcal{L}_{\text{hrchyLD}}$ 35.58 52.85 34.25 50.96 DA only distillation 36.42 53.32 34.05 50.83 only teacher training 37.28 53.65 32.71 49.72 not at all 36.43 53.68 33.54 50.93 NECHO v2 Full 37.02 54.26 34.24 51.01
+Criteria Components (0.2, 0.2, 0.2) (0.5, 0.2, 0.8) top-10 top-20 top-10 top-20 KD w/o $\mathcal{L}_{\text{MWCD}}$ 37.10 54.10 34.37 50.79 w/o $\mathcal{L}_{\text{TR2D}}$ 36.9 53.95 33.15 50.58 w/o $\mathcal{L}_{\text{MAGD}}$ 36.01 53.27 32.91 49.87 w/o $\mathcal{L}_{\text{hrchyLD}}$ 35.58 52.85 34.25 50.96 DA Only During Distillation 36.42 53.32 34.05 50.83 Only During Teacher Training 37.28 53.65 32.71 49.72 Not For Both 36.43 53.68 33.54 50.93 NECHO v2 Full 37.02 54.26 34.24 51.01
 Table 3: Ablation Studies on MIMIC-III Data.
 To evaluate our proposed components, we conduct ablation studies on MIMIC-III data, as detailed in Table 3.
-We report two scenarios: a balanced missing ratio of 0.2, and an imbalanced ratios of (0.5, 0.8, 0.2), representing two extremes where diagnosis codes representation is either highly dominant or minimal.
+We report two scenarios: a balanced missing ratio of 0.2, and an imbalanced ratios of (0.5, 0.8, 0.2), representing two extremes where diagnosis codes representations are either highly dominant or minimal.
 We first assess the effectiveness of KD.
-Whilst $\mathcal{L}_{\text{MWCD}}$ occasionally performs better without it, its consistent use generally enhances performance.
-The absence of $\mathcal{L}_{\text{TR2D}}$ and $\mathcal{L}_{\text{MAGD}}$ during distillation significantly deteriorates the performance, highlighting the importance of intermediate supervision.
-Additionally, $\mathcal{L}_{\text{hrchyLD}}$ is beneficial in performance.
-The quantitative results validate the importance of all components in systematic KD to align student’s semantic knowledge to that of the teacher.
-We also evaluate the efficacy of DA in three scenarios: only distillation, only teacher training, and not at all.
-Overall performance significantly improves, highlighting the significance of the proposed random single data point erasing under missing visit information.
-This enhances the teacher’s robustness against missingness during training and minimises data distribution discrepancies during distillation, resulting in a student model that is highly resilient to uncertain data incompleteness.
+Whilst NECHO v2 occasionally performs better without $\mathcal{L}_{\text{MWCD}}$ and $\mathcal{L}_{\text{MWHD}}$, their consistent use generally enhances performance.
+The absence of $\mathcal{L}_{\text{TR2D}}$ and $\mathcal{L}_{\text{MAGD}}$ during distillation significantly deteriorates the performance, highlighting the importance of intermediate representation propagation.
+Additionally, $\mathcal{L}_{\text{hrchyLD}}$ is beneficial.
+These validate the importance of all components in our systematic KD pipeline to align the student’s semantic knowledge to that of the teacher.
+We also evaluate the efficacy of data erasing against three scenarios: only during distillation, only during teacher training, and not for both.
+Overall performance considerably improves, highlighting the significance of the proposed curriculum random data erasing under missing visit information.
+This enhances the teacher’s robustness against missingness during training and minimises data distribution discrepancies during distillation, resulting in the student model that is highly resilient to uncertain data incompleteness.
 3.2.3 Comparative Studies
-Criteria Components (0.2, 0.2, 0.2) (0.5, 0.2, 0.8) top-10 top-20 top-10 top-20 Pairing Original $\to$ Original 36.79 53.11 32.68 49.53 Original $\to$ Modified w/ MAG 36.44 52.96 34.66 50.99 $\mathcal{L}_{\text{TR2D}}$ Not Random 36.91 53.56 32.83 50.40 NECHO v2 Proposed 37.02 54.26 34.24 51.01
+Criteria Components (0.2, 0.2, 0.2) (0.5, 0.2, 0.8) top-10 top-20 top-10 top-20 Pairing Original $\to$ Original 36.79 53.11 32.68 49.53 Original $\to$ Modified for Student 36.44 52.96 34.66 50.99 $\mathcal{L}_{\text{TR2D}}$ Not Random 36.91 53.56 32.83 50.40 NECHO v2 Proposed 37.02 54.26 34.24 51.01
 Table 4: Comparative Studies on MIMIC-III Data.
-As shown in Table 4, we compare NECHO v2 with various combinations of teacher and student (original to original, original to modified with MAG) and transformer full distillation under the same settings with the ablation studies.
-Our proposed methodologies achieves the best overall performance, highlighting the importance of: 1) carefully pairing teacher and student to address varying representation dominance and minimise architectural heterogeneity, and 2) introducing randomness into KD to prevent overfitting and foster effective knowledge transfer.
+Under the same settings as the ablation studies, we compare our NECHO v2 with different teacher-student combinations (original to original, original to modified for student) and transformer not random distillation.
+Our proposed methodologies achieve the best overall performance, underscoring the importance of: 1) carefully pairing teacher and student to address shifting representation dominance and minimise architectural heterogeneity, and 2) incorporating randomness into KD to prevent overfitting.
+We provide the corresponding result to Table 4.
 4 Conclusion
 We tackle uncertain missing sequences for robust multimodal SDP with the proposed NECHO v2.
-Extensive experiments show the stable superiority of NECHO v2 over the existing methods.
-The code will be released soon.
+With modified NECHO that dynamically adjusts dominant representations under varying missingness, we design a curriculum data erasing guided systematic KD pipeline that enables the student to effectively imitate the teacher.
+Extensive experiments on MIMIC-III data show the effectiveness of our approach over the existing methodologies.
+To foster future research, we release code at: https://www.github.com/heejkoo9/NECHOv2.

### v4

@@
-Overcoming Uncertain Incompleteness for Robust Multimodal Sequential Diagnosis Prediction via Curriculum Data Erasing Guided Distillation
+Overcoming Uncertain Incompleteness for Robust Multimodal Sequential Diagnosis Prediction via Curriculum Data Erasing Guided Knowledge Distillation
 Abstract
@@
 Encountering such situations presents a formidable challenge to healthcare analytics.
-Missing data, a common issue in reality, exacerbates model performance.
+Missing data, a common issue in reality, exacerbates model performance [5].
 Basic approaches, such as imputation by mean or excluding incomplete data instances, often fail to preserve true data distribution and result in information loss [6].
@@
 $\mathcal{L}_{\text{MWCD}}^{\text{student $\rightarrow$ teacher},m}=-\log\frac{\exp(\langle\hat{R}_{t}^{\text{student},m,i},\hat{R}_{t}^{\text{teacher},m,i}\rangle/\tau)}{\sum_{k=1}^{K}\exp(\langle\hat{R}_{t}^{\text{student},m,i},\hat{R}_{t}^{\text{teacher},m,k}\rangle/\tau)},$ (2)
-$\mathcal{L}_{\text{MWCD}}=\sum_{m\{D,N,C\}}\{\alpha\mathcal{L}_{\text{MWCD}}^{\text{\text{teacher}, $\rightarrow$ student},m}+(1-\alpha)\mathcal{L}_{\text{MWCD}}^{\text{student $\rightarrow$ \text{teacher}, },m}\}$ (3)
+$\mathcal{L}_{\text{MWCD}}=\sum_{m\in\{D,N,C\}}\{\alpha\mathcal{L}_{\text{MWCD}}^{\text{\text{teacher}, $\rightarrow$ student},m}+(1-\alpha)\mathcal{L}_{\text{MWCD}}^{\text{student $\rightarrow$ \text{teacher}, },m}\}$ (3)
 where $\langle$,$\rangle$ , ⟩ denotes cosine similarity and the temperature $\tau\in\mathbb{R}^{+}$ is a parameter that controls the distribution concentration and the gradient of the Softmax function.
@@
 Firstly, the teacher is trained using curriculum learning guided random data erasing, starting with easier samples and gradually progressing to more difficult ones.
-Each modality is assigned a missing probability of 0.0 or 0.1 with equal probability until specific epoch $e_{1}$, after which the probability of 0.2 is added.
-Thereafter, during the distillation, complete data representations from the teacher trained in the previous manner are migrated until epoch $e_{2}$, after which training continues with either no missing data or a 0.1 missing ratio.
+All modalities are assigned a missing probability of 0.0 or 0.1 with equal probability until specific epoch $e_{1}$, after which the probability of 0.2 is added.
+Thereafter, during the distillation, complete data representations from the teacher trained in the previous manner are migrated until epoch $e_{2}$, after which training continues with either no missing data or a 0.1 missing ratio to each modality.
 This strategy improves robustness of the teacher against missing data during training and reduces data distribution discrepancies during distillation, leading to an improved representation transmission.
@@
 For instance, if the total number of codes is 3882, the missing value is assigned as 3883.
-We also replace missing tokens in clinical notes with the UNK token [9].
+We also replace missing tokens in clinical notes with UNK token [10].
 Criteria MIMIC-III Count General # of Patients 5551 # of Unique Codes 3882 # of Category Codes 126 # of Typing Codes 17 Visit # of Visits 14568 Avg / Max # Visit per Patient 3.37 / 33 Avg / Max # Unique Codes per Visit 13.29 / 39 Avg / Max # Category Codes per Visit 11.46 / 34 Avg / Max # Typing Codes per Visit 6.71 / 15